A weightless bar 2 metres long is suspended from the ceiling at a point
which is 0.5 metres in from one end. Suspended from the same end is a
mass of 110 kg. Find the mass which must be suspended from a point 0.3
metres in from the other end of the bar so that the bar will remain
horizontal.
1
Expert's answer
2020-09-24T11:08:36-0400

The bar will be horizontal if the moments opposing each other are equal:


"M_1+M_2=0,\\
m_1gd_1-m_2gd_2=0,\\
m_2=m_1\frac{d_1}{d_2}=110\cdot\frac{0.5}{0.3}=183.33\text{ kg}."